# Magneto: Combining Small and Large Language Models for Schema Matching

### v1

@@ 5. The GDC Benchmark
A key problem in evaluating schema-matching algorithms is the availability of benchmarks that cover a wide variety of challenges present in real data. While there have been efforts to address this issue, benchmarks are often synthetically fabricated from real data (Koutras et al., 2021; Das et al., [n.d.]), which can introduce biases. Moreover, as we show in Section 6 (and as seen in (Koutras et al., 2021)), publicly available benchmarks based on real data such as Magellan (Das et al., [n.d.]) and WikiData (Koutras et al., 2021) are quickly becoming “saturated” since many algorithms attain near-perfect performance and leave small room for algorithmic improvements. This makes it difficult to extract useful insights about the strengths of different algorithms.
@@
To address this problem, we developed a new benchmark dataset based on the real data harmonization scenario described in Example 1.1. We collaborated with experts from a medical school team to design a benchmark that reflects the challenges they face when working with biomedical data. We obtained datasets from ten studies related to tumor analysis (Cao et al., 2021; Clark et al., 2019; Dou et al., 2020; Gillette et al., 2020; McDermott et al., 2020; Huang et al., 2021; Krug et al., 2020; Satpathy et al., 2021; Vasaikar et al., 2019; Wang et al., 2021), and, with the help of the experts, manually aligned and matched these datasets to the Genomics Data Commons (GDC) standard (National Cancer Institute, 2024a).
@@
The GDC is a program of the US National Institutes of Health responsible for receiving, processing, and distributing genomic, clinical, and biospecimen data from cancer research initiatives. Its standard dictionary describes data using a graph model that includes names and descriptions for nodes and attributes, as well as acceptable values for some attributes. Still, for compatibility with existing solutions, we transformed the model to a relational schema that contains only column names and domain information (i.e., we disregard column descriptions). We created a simplified table reflecting the GDC format, the “target” table, listing domain values for each column without repetition. In a future version, it is possible to further extend the benchmark to include multiple tables, and column and value descriptions.
@@
Table 1 illustrates a few samples of the data published by Dou et al. (Dou et al., 2020) alongside their corresponding GDC format. As we can observe, matching clinical data with the GDC standard poses several challenges, including terminology mismatches and data format variations. The benchmark includes 10 pairs of input-target tables. The number of columns in the source tables ranges from 16 to 179, and the number of rows ranges from 93 to 225. Our simplified GDC target schema comprises a single table with 736 columns. While some columns have a small number of distinct values (e.g., binary yes/no attributes), some contain up to 4478 distinct values. The ground truth was manually curated by multiple annotators, who used a mix of manual and automated methods for identifying possible candidate matches (e.g., GDC search tools (National Cancer Institute, 2024b) and bdi-kit (The bdi-kit contributors, 2024)). Given that the correctness of some matches is very challenging to determine even for bioinformatics experts (e.g., it may require reading the original papers or asking data producers), the final match decisions were made by consensus based on what users would expect from an algorithm given the limited context.
@@ 6.1. Experimental Setup
Datasets. We evaluated Magneto using the six datasets detailed in Table 3, which we group into two categories: (1) real – human-curated and capture column matches seen in the real world, and (2) synthetic –systematically generated to capture structural variations and diverse match types. The GDC Benchmark (Section 5) reflects challenges in biomedical data integration scenario. The other five datasets are from the Valentine schema matching benchmark (Koutras et al., 2021) and have been used to evaluate recent schema matching solutions (Tu et al., 2023; Du et al., 2024).
@@
Dataset #Table Pairs #Cols #Rows Type
GDC 10 16–733 93–4.5k Real
Magellan 7 3–7 0.9k–131k Real
WikiData 4 13–20 5.4k–10.8k Real
Open Data 180 26–51 11.6k–23k Synthetic
ChEMBL 180 12–23 7.5k–15k Synthetic
TPC-DI 180 11–22 7.5k–15k Synthetic
Baselines. We compare Magneto against several approaches, ranging from traditional string similarity-based methods to sophisticated model-driven techniques. For traditional approaches, we use:
@@
We evaluated the Jaccard-Levenshtein baseline proposed in (Koutras et al., 2021) and the Cupid algorithm (Madhavan et al., 2001). However, since they consistently performed significantly worse than the other methods, we exclude them from the reporting for clarity. We used the implementations available in the Valentine repository these methods (Koutras et al., 2021).
We also used recent approaches that use LMs: ISResMat which leverages contrastive learning and embeddings from SLMs (Du et al., 2024), and Unicorn, a supervised and general approach using language models for data encoding and trained for data integration tasks (Tu et al., 2023). We used the implementation from the authors for ISResMat. For Unicorn, we used the pre-trained model and implementation provided by the authors. The match scores were derived from the model’s predicted match probabilities. It is important to note that Unicorn requires external training data for optimal performance. The zero-shot version performed significantly worse than other baselines and was excluded from our evaluation. We report Unicorn numbers for the algorithm in Valentine datasets for completeness, but note that its model was trained on the fabricated datasets from Valentine (Koutras et al., 2021), so data leakage can influence the reported results.
@@
Execution platform. The experiments were ran on a server running Red Hat Enterprise Linux version 9.2. We used a single node consisting of 8 cores and 64 GB of memory for CPU-based tasks. The GPU-intensive experiments used an NVIDIA A100 GPU. For Magneto variations, Unicorn, and IsResMat, we used a GPU to run all experiments, except the scalability experiment described in Section 6.3, conducted on a CPU for consistency with other methods.
@@ 6.2. End-to-End Accuracy
Performance on Valentine. The accuracy results for the Valentine benchmark are shown in Figure 7. Note that the different methods attain much higher accuracies for these datasets than for GDC. The Valentine datasets are derived either from open data sources or synthetic repositories, and they are also meticulously curated. In contrast, the GDC benchmark includes real scientific datasets and features a diverse set of column names and values, which, as these results confirm, is particularly challenging for SM solutions.
Figure 7. Accuracy on Valentine datasets. Magneto variants attain high accuracy (even without fine-tuning or data-specific training) compared to traditional, model-based, and supervised approaches.
@@
The high accuracy of the zero-shot configurations of Magneto leaves little room for improvement. Specifically, fine-tuning faces challenges within the Valentine datasets, which predominantly contains lexical rather than semantic matches. For instance, the OpenData dataset includes matches like Gender to Ge and Employer to Em, which are uncommon in practical applications. Despite these obstacles, our strategy of fine-tuning combined with LLM-based re-ranking generally enhances performance. Additionally, we demonstrate in Section 5 that combining simple synthetic data generation techniques with LLM-derived data leads to better performance in this context.
@@ 6.3. Scalability Assessment
We compare the runtime of Magneto to those of the top three baselines in accuracy: Unicorn, Coma++, and ISResMat. Since Coma++ does not support GPU execution, we ran this experiment in CPU mode to ensure a fair comparison. We report only the resuls of Magneto-zs-bp and Magneto-zs-llm since their runtime is similar to Magneto-ft-bp and Magneto-ft-llm, respectively.
@@
Figure 8. Runtime analysis using a log scale on the y-axis. Magneto shows good scalability in scenarios that involve a large number of columns and rows. Magneto-zs-bp is often significantly faster than the other baselines, while Magneto-zs-llm maintains a consistent runtime, completing tasks within minutes, even for datasets with hundreds of columns. Note that ISResMat and Unicorn fail to complete execution within the time limit for GDC after 100 columns.
The performance of Magneto-zs-bp and Magneto-zs-bp is stable as the table size increases; with Magneto-zs-bp showing minimal runtime growth and Magneto-zs-llm demonstrating runtime stability despite its complexity. For GDC, Magneto-zs-bp shows runtimes ranging from 11 to 33 seconds. Such a low increase in runtime reflects its efficient design, bounded primarily by the efficient embedding computations. As expected, Magneto-zs-llm incurs higher runtimes due to the LLM requests. Still, it shows stability for increasing the number of columns with a runtime starting at 545 and peaking at 589 seconds since the amount of data sent does not change as we increase the number of columns in the target. The runtimes of Magneto variations are even lower for OpenData: the dataset has fewer columns but many more rows, which Magneto compensates through sampling.
Coma++ shows low runtimes for a small number of columns, but its performance decreases as the number of columns grows. IsResMat and Unicorn exhibit significant scalability challenges, with their runtime growing substantially with the number of columns. IsResMat was not able to complete the execution for GDC, not even for the lowest number of columns, and its runtime for OpenData was orders of magnitude higher than the other methods. Unicorn could only process the initial 100 columns for GDC.
@@ 6.4. Ablation of Magneto Components
In zero-shot settings, $\mathcal{S}_{\texttt{repeat}}$ consistently outperforms other methods, demonstrating the effectiveness of emphasizing column header when training data is absent. $\mathcal{S}_{\texttt{verbose}}$ is more effective in fine-tuned scenarios, demonstrating the benefits of integrating semantic details when domain-specific training is applied. The Header Only approach shows the lowest performance in all datasets, confirming that column headers alone are insufficient for effective schema matching in complex scenarios.
The results further support our assumption that that fine-tuning can greatly improve the performance gains on the GDC dataset. This improvement is likely related to the requirement for external knowledge, which we capture with the LLM-derived training data.
@@
Value Sampling. We compare the effectiveness our proposed Priority Sampling (Section 3.1) against the following sampling methods: Coordinated, a variation of Priority that excludes frequency weights; Weighted and Frequency, variations of Prioriry without coordination but with value frequencies – Frequency uses the most frequent values and Weighted uses weighted sampling based on these frequencies; Random uses basic random sampling.
We used GDC and the three synthetic datasets from Valentine; we exclude Wikidata and Magellan since as all methods already perform well on them. A the results in Table 4 show, Magneto-zs-bp with $\mathcal{S}_{\texttt{repeat}}$, the best zero-shot setting configuration, Priority, generally outperforms other methods in both Recall@GT and MRR, and sometimes is a close second. Priority Sampling prioritizes frequently occurring values and enhances the likelihood of sampling similar values across columns, a beneficial feature for schema matching. The top three techniques—including Priority and its ablations—perform comparably well, also representing viable choices.
Synthetic Data Generation. We evaluate our LLM-powered synthetic data generation for SLM fine-tuning against two other methods: simple augmentation, which incorporates row shuffling, sampling, and column name perturbation, and mixed, which combines these methods. We used the optimal settings from serialization studies, $\mathcal{S}_{\texttt{repeat}}$, on Magneto-ft-bp and Magneto-ft-llm. We assess MRR for both methods and Recall@GT for Magneto-ft-bp, as Recall@GT for Magneto-ft-llm is largely influenced by the reranker instead of the fine-tuned candidate retriever. This analysis focuses on the four datasets that have the potential for improvement.
The results in Table 5 indicate that LLM-based and mixed data generation significantly outperform simple perturbations on the challenging GDC dataset. However, on the fabricated Valentine datasets, synthetic training data generated by LLMs shows limited impact due to the datasets nature, which feature uncommon and semantically sparse matches less amenable to LLM-derived enhancements. Nevertheless, the marked performance differential on the GDC dataset highlights the effectiveness of our fine-tuning strategy for real data.
@@
Table 5. Ablation of data generation techniques. These results show that LLM-powered data generation is effective, and better performance can be achieved by combining multiple techniques.
@@
LLM-based 0.866±0.083 0.960±0.089 0.939±0.080 0.971±0.079
Mixed 0.830±0.077 0.949±0.109 0.927±0.096 0.965±0.096
Simple 0.798±0.116 0.955±0.097 0.917±0.105 0.969±0.081
@@
Match Reranker We performed two experiments to evaluate the match reranker component. The first examines the impact of using the bipartite strategy (Section 2.2). Figure 10 shows the absolute increase/decrease in accuracy from Magneto-zs, which directly returns the retriever matches, to Magneto-zs-bp, which applies the bipartite reranking to the matches produced by the retriever. Note that the bipartite approach can significantly improve Recall@sGT. All datasets benefit from the technique at some point, except Magellan, whose scores are already (very) high. The improvements can reach nearly 0.6 in absolute terms. The mean improvement varies across datasets; average recall scores across datasets other than Magellan range from 0.03 (GDC) to 0.17 (OpenData). The improvements for MRR are lower but still sizeable for ChemL and Open Data. While reranking with bipartite generally works, we observed that it can occasionally lead to decreased MRR and Recall@GT.
Our final experiment evaluates: (1) how the LLM-based reranker improves upon the traditional bipartite approach, and (2) how its performance varies as we increase the number of match candidates sent to the LLM. We focus on the GDC dataset, which has a higher column count. For this experiment, we used the zero-shot settings for the retriever. Figure 11 shows accuracy and runtime for different numbers of candidates ($k$).
@@
Figure 11. Ablation of number of candidates sent to LLM-based reranker on GDC.
The LLM-based re-ranking approach improves over the bipartite. For example, at $k=5$, we observe a 12.3% absolute improvement in Recall@GT over the bipartite baseline, rising from 0.404 to 0.454. Mean Reciprocal Rank (MRR) demonstrates even more promising gains, from 0.551 to 0.605—a 9.8% increase. While MRR continues to improve with more columns, processing all target table columns becomes computationally infeasible for large schemas. That is why we argue for combining a retriever with an LLM-based re-ranker for schema matching. Note that candidates that are not forwarded to the reranker are assigned lower scores, this leads to a decrease in Recall@GT since it considers all pairs based on these scores.
As the number of columns increases, computational costs grow. The runtime reflects the latency of API-based LLM requests. However, we can expect similar trends with alternative LLM approaches, as prompts with increasing length become more challenging for LLMs (Wan et al., 2024). Still, the runtime remains practical and competitive for usable values of $k$, as demonstrated in our scalability experiment. We used $k=20$ for our other experiments, as it balances good MRR, Recall@GT, and practical runtime across GDC and other datasets.

### v2

@@ 5. The GDC benchmark
A major challenge in evaluating schema-matching algorithms is the lack of benchmarks reflecting the diversity of real-world data. Benchmarks are often synthetically fabricated from real data (Koutras et al., 2021; Das et al., [n.d.]), which can introduce biases. Moreover, as show in Section 6 (and in (Koutras et al., 2021)), publicly available benchmarks based on real data such as Magellan (Das et al., [n.d.]) and WikiData (Koutras et al., 2021) are quickly becoming “saturated” since many algorithms attain near-perfect performance and leave small room for algorithmic improvements. This makes it difficult to extract useful insights about the strengths of different algorithms.
@@
To address this problem, we built a new benchmark dataset (Santos et al., 2025) based on the real data harmonization scenario described in Example 1.1. We collaborated with biomedical researchers to design a benchmark that reflects the challenges they face when working with biomedical data. We obtained datasets from ten studies related to tumor analysis (Cao et al., 2021; Clark et al., 2019; Dou et al., 2020; Gillette et al., 2020; McDermott et al., 2020; Huang et al., 2021; Krug et al., 2020; Satpathy et al., 2021; Vasaikar et al., 2019; Wang et al., 2021), and, with the experts’ help, manually aligned and matched these datasets to the Genomics Data Commons (GDC) standard (National Cancer Institute, 2024a).
@@
The GDC is a program of the US National Institutes of Health responsible for handling genomic, clinical, and biospecimen data from cancer research initiatives. Its standard dictionary describes data using a graph model that includes names and descriptions for nodes and attributes and acceptable values for some attributes. To be compatible with existing schema matching solutions, we transformed the model to a relational schema that contains only column names and domain information (i.e., we disregard column descriptions). We created a simplified table reflecting the GDC format, the “target” table, listing domain values for each column without repetition.
@@
Table 1 illustrates a few samples of the data published by Dou et al. (Dou et al., 2020) alongside their corresponding GDC format. As we can observe, matching clinical data with the GDC standard poses several challenges, including terminology mismatches and data format variations. The benchmark includes 10 pairs of source-target tables. The number of columns in the source tables ranges from 16 to 179, and the number of rows ranges from 93 to 225. Our simplified GDC target schema comprises a single table with 736 columns. While some columns have a small number of distinct values (e.g., binary yes/no attributes), some contain up to 4478 distinct values. The ground truth was manually curated by multiple annotators, who used a mix of manual and automated methods to identify possible candidate matches (e.g., GDC search tools (National Cancer Institute, 2024b) and bdi-kit (The bdi-kit contributors, 2024)). Given that the correctness of some matches is challenging to determine even for bioinformatics experts (e.g., it may require reading the original papers or asking data producers), the final match decisions were made by consensus based on what users would expect from an algorithm given the limited context.
@@ 6.1. Experimental Setup
Datasets. We evaluated Magneto on six datasets (Table 3), grouped as: (1) Human-Curated-reflecting real-world matches, and (2) Fabricated—systematically generated to capture structural variations and diverse match types. The GDC benchmark (Section 5) reflects challenges in biomedical data integration. The other five datasets are from the Valentine schema matching benchmark (Koutras et al., 2021) and have been used to evaluate recent schema matching solutions (Tu et al., 2023; Du et al., 2024).
@@
Dataset #Table Pairs #Cols #Rows Match Type
GDC 10 16–736 93–4.5k Human-Curated
Magellan 7 3–7 0.9k–131k Human-Curated
WikiData 4 13–20 5.4k–10.8k Human-Curated
Open Data 180 26–51 11.6k–23k Fabricated
ChEMBL 180 12–23 7.5k–15k Fabricated
TPC-DI 180 11–22 7.5k–15k Fabricated
Baselines. We compare Magneto against several approaches, ranging from traditional string-similarity-based methods to sophisticated model-driven techniques. For traditional approaches, we use:
@@
We evaluated the Jaccard-Levenshtein baseline proposed in (Koutras et al., 2021) and the Cupid algorithm (Madhavan et al., 2001). However, both methods consistently performed significantly worse than other approaches across all experiments, so we excluded them from the plots for clarity. We used the implementations available in the Valentine repository for all traditional approaches (Koutras et al., 2021).
We also compared Magneto against recent approaches that use LMs: ISResMat which leverages contrastive learning and embeddings from SLMs (Du et al., 2024), and Unicorn, a supervised and general approach using language models for data encoding and trained for data integration tasks (Tu et al., 2023). We used the implementation from the authors for ISResMat. For Unicorn, we used the pre-trained model and implementation provided by the authors. The match scores were derived from the model’s predicted match probabilities. It is important to note that Unicorn requires external training data for optimal performance. The zero-shot version performed significantly worse than other baselines and was excluded from our evaluation. We report Unicorn numbers for the algorithm in Valentine datasets for completeness, but note that its model was trained on the fabricated datasets from Valentine (Koutras et al., 2021), so data leakage can influence the reported results.
@@
Execution platform. The experiments were run on a server running Red Hat Enterprise Linux version 9.2. We used a single node consisting of 8 cores and 64 GB of memory for CPU-based tasks. The GPU-intensive experiments used an NVIDIA A100 GPU. For Magneto variations, Unicorn, and IsResMat, we used a GPU to run all experiments, except the scalability experiment described in Section 6.3, conducted on a CPU for consistency with other methods.
@@ 6.2. End-to-End Accuracy
Performance on Valentine. The accuracy results for the Valentine benchmark are shown in Figure 7. Note that the different methods attain much higher accuracies for these datasets than for GDC. The Valentine datasets are derived either from open data sources or synthetic repositories, and they are also meticulously curated. In contrast, the GDC benchmark includes real scientific datasets and features a diverse set of column names and values, which, as these results confirm, is particularly challenging for schema matching solutions.
Figure 7. Accuracy on Valentine datasets. Magneto variants attain high accuracy (even without fine-tuning) compared to traditional, model-based, and supervised approaches.
@@
The high accuracy of the zero-shot configurations of Magneto leaves little room for improvement. Specifically, fine-tuning faces challenges within the Valentine datasets, which predominantly contain lexical rather than semantic matches. For instance, the OpenData dataset includes matches like Gender to Ge and Employer to Em, which are uncommon in practical applications. Despite these obstacles, our LLM-based fine-tuning and re-ranking generally enhance performance. Additionally, we demonstrate in Section 6 that combining simple synthetic data generation techniques with LLM-derived data leads to better performance in this context.
@@ 6.3. Scalability Assessment
We compare the runtime of Magneto to those of the top three baselines in accuracy: Unicorn, Coma++, and ISResMat. Since Coma++ does not support GPU execution, we ran this experiment in CPU mode to ensure a fair comparison. We report only the results of Magneto-zs-bp and Magneto-zs-llm since their runtime is similar to Magneto-ft-bp and Magneto-ft-llm, respectively.
@@
Figure 8. Runtime analysis (log-scale y-axis). Magneto scales well with large datasets: Magneto-zs-bp is often much faster than baselines, while Magneto-zs-llm maintains consistent runtimes, finishing large tasks within minutes. ISResMat and Unicorn fail to complete GDC after 100 columns.
Magneto-zs-bp and Magneto-zs-llm remain stable with increasing table size: the runtime for Magneto-zs-bp grows slightly and Magneto-zs-llm maintains a stable runtime despite its complexity. For GDC, Magneto-zs-bp shows runtimes ranging from 11–33 seconds. Such a low increase in runtime reflects its efficient design, bounded primarily by the embedding computations. As expected, Magneto-zs-llm incurs higher runtimes due to the LLM requests (545–589 seconds), but it shows stability as the number of columns increases since the amount of data sent does not change. The runtimes of Magneto variations are even lower for OpenData: the dataset has fewer columns but many more rows, which Magneto compensates through sampling.
Coma++ shows low runtimes for a small number of columns, but its performance decreases as the number of columns grows. IsResMat and Unicorn exhibit significant scalability challenges, as their runtime grows substantially with the number of columns. IsResMat was not able to complete the execution for GDC, not even for the lowest number of columns, and its runtime for OpenData was orders of magnitude higher than the other methods. Unicorn could only process the initial 100 columns for GDC.
@@ 6.4. Ablation of Magneto Components
In zero-shot settings, $\mathcal{S}_{\texttt{repeat}}$ consistently outperforms other methods, demonstrating the effectiveness of emphasizing the column header and raising its cumulative attention in the zero-shot regime. While $\mathcal{S}_{\texttt{verbose}}$ under-performs in zero-shot settings, it excels after fine-tuning because the prefixes act as learnable anchors that guide the model to separate and re-weight heterogeneous components (name, type, values) during representation learning, highlighting the benefits of integrating semantic details when domain-specific training is applied. The Header only-ZS approach shows the lowest performance in all datasets, confirming that column headers alone are insufficient in complex scenarios.
The results further support our assumption that fine-tuning can greatly improve the performance gains on the GDC dataset. This improvement is likely related to the requirement for external knowledge, which we capture with the LLM-derived training data.
@@
Value Sampling. We compare the effectiveness our proposed Priority Sampling (Section 3.1) against the following sampling methods: Coordinated, a variation of Priority that excludes frequency weights; Weighted and Frequency, variations of Priority without coordination but with value frequencies – Frequency uses the most frequent values and Weighted uses weighted sampling based on these frequencies; Random uses basic random sampling.
We used GDC and the three synthetic datasets from Valentine; we exclude Wikidata and Magellan since as all methods already perform well on them. As the results in Table 4 show, Magneto-zs-bp with $\mathcal{S}_{\texttt{repeat}}$, the best zero-shot setting configuration, Priority, generally outperforms other methods in both Recall@GT and MRR, and sometimes is a close second. Priority Sampling prioritizes frequently occurring values and enhances the likelihood of sampling similar values across columns, a beneficial feature for schema matching. The top three techniques—including Priority and its ablations—perform comparably well, also representing viable choices.
Synthetic Data Generation. We evaluate our llm-augfor SLM fine-tuning against two other methods: struct-aug, which incorporates row shuffling, sampling, and column name perturbation, and mixed-aug, which combines these methods. We used the optimal settings from serialization studies, $\mathcal{S}_{\texttt{repeat}}$, on Magneto-ft-bp and Magneto-ft-llm. We assess MRR for both methods and Recall@GT for Magneto-ft-bp, as Recall@GT for Magneto-ft-llm is largely influenced by the reranker instead of the fine-tuned candidate retriever. This analysis focuses on the four datasets that have the potential for improvement. Table 5 shows that llm-augand mixed-aug outperform simple perturbations on the GDC dataset. However, on Valentine datasets, LLM-generated data has limited impact due to sparse, uncommon matches. Still, the strong GDC results highlight the effectiveness of our fine-tuning for real data.
@@
Table 5. Ablation of data generation techniques. LLM-powered data generation is effective, and better performance can be achieved by combining multiple techniques.
@@
llm-aug 0.866±0.083 0.960±0.089 0.939±0.080 0.971±0.079
mixed-aug 0.830±0.077 0.949±0.109 0.927±0.096 0.965±0.096
struct-aug 0.798±0.116 0.955±0.097 0.917±0.105 0.969±0.081
Model Ablations. We extended our ablation studies to include RoBERTa (Liu, 2019) and E5 (Wang et al., 2022) (SLMs), and LLaMA3.3-70B (Dubey et al., 2024) (LLM). Table 6 shows the results. Regarding SLMs the default MPNet outperforms other SLMs in most variations of Magneto. In general, the serialization and score-based prompt proposed are robust. While $\mathcal{S}_{\texttt{repeat}}$ dominates most zero-shot settings, $\mathcal{S}_{\texttt{verbose}}$ and $\mathcal{S}_{\texttt{default}}$ yield even higher gains after fine-tuning, which is consistent with our findings in Figure 9. The tested models benefit from the same prompting framework, which confirms that Magneto performs consistently using different models with a robust cost-accuracy tradeoff. For LLMs, LLaMA3.3-70B surpasses GPT-4o-mini (MRR=0.837 vs. 0.815), confirming that larger LLMs enhance reranking. Combining LLaMA3.3-70B with fine-tuned MPNet achieves the best overall performance (MRR=0.860), reinforcing the value of having a pipeline that performs retrieval and reranking.
Table 6. Model ablation comparing three SLMs and two LLMs with MRR///Recall@GT metrics. Bold indicates best serialization per model;     shows best zero-shot model and     best fine-tuned model.
Model zs-bp ft-bp zs-gpt4o-mini ft-gpt4o-mini zs-llama3.3-70b ft-llama3.3-70b
MPNet ($\mathcal{S}_{\texttt{default}}$) 0.656$\pm$0.082 / 0.357$\pm$0.093 0.740$\pm$0.089 / 0.400$\pm$0.094 0.775$\pm$0.102 / 0.412$\pm$0.115 0.846$\pm$0.104 / 0.537$\pm$0.162 0.772$\pm$0.101 / 0.339$\pm$0.155 0.860$\pm$0.088 / 0.423$\pm$0.170
MPNet ($\mathcal{S}_{\texttt{verbose}}$) 0.627$\pm$0.119 / 0.341$\pm$0.114 0.761$\pm$0.071 / 0.438$\pm$0.085 0.758$\pm$0.118 / 0.469$\pm$0.131 0.830$\pm$0.766 / 0.479$\pm$0.154 0.776$\pm$0.098 / 0.323$\pm$0.153 0.838$\pm$0.121 / 0.446$\pm$0.185
MPNet ($\mathcal{S}_{\texttt{repeat}}$) 0.731$\pm$0.086 / 0.375$\pm$0.108 0.701$\pm$0.110 / 0.172$\pm$0.103 0.808$\pm$0.095 / 0.430$\pm$0.131 0.820$\pm$0.117 / 0.524$\pm$0.104 0.828$\pm$0.096 / 0.358$\pm$0.117 0.819$\pm$0.102 / 0.419$\pm$0.138
RoBERTa ($\mathcal{S}_{\texttt{default}}$) 0.631$\pm$0.105 / 0.336$\pm$0.081 0.703$\pm$0.120 / 0.378$\pm$0.107 0.734$\pm$0.119 / 0.381$\pm$0.088 0.784$\pm$0.131 / 0.490$\pm$0.149 0.755$\pm$0.098 / 0.292$\pm$0.119 0.815$\pm$0.121 / 0.432$\pm$0.164
RoBERTa ($\mathcal{S}_{\texttt{verbose}}$) 0.621$\pm$0.063 / 0.350$\pm$0.083 0.692$\pm$0.114 / 0.367$\pm$0.112 0.743$\pm$0.097 / 0.400$\pm$0.099 0.821$\pm$0.092 / 0.511$\pm$0.160 0.751$\pm$0.075 / 0.324$\pm$0.140 0.854$\pm$0.101 / 0.424$\pm$0.147
RoBERTa ($\mathcal{S}_{\texttt{repeat}}$) 0.710$\pm$0.103 / 0.373$\pm$0.097 0.681$\pm$0.096 / 0.368$\pm$0.111 0.794$\pm$0.097 / 0.434$\pm$0.173 0.774$\pm$0.101 / 0.504$\pm$0.161 0.814$\pm$0.088 / 0.398$\pm$0.161 0.810$\pm$0.090 / 0.428$\pm$0.165
E5 ($\mathcal{S}_{\texttt{default}}$) 0.623$\pm$0.104 / 0.329$\pm$0.105 0.729$\pm$0.081 / 0.406$\pm$0.094 0.738$\pm$0.144 / 0.356$\pm$0.144 0.832$\pm$0.080 / 0.480$\pm$0.117 0.756$\pm$0.120 / 0.405$\pm$0.090 0.857$\pm$0.076 / 0.388$\pm$0.112
E5 ($\mathcal{S}_{\texttt{verbose}}$) 0.602$\pm$0.094 / 0.322$\pm$0.102 0.692$\pm$0.114 / 0.367$\pm$0.112 0.745$\pm$0.131 / 0.382$\pm$0.135 0.821$\pm$0.092 / 0.511$\pm$0.160 0.746$\pm$0.100 / 0.327$\pm$0.100 0.838$\pm$0.121 / 0.446$\pm$0.185
E5 ($\mathcal{S}_{\texttt{repeat}}$) 0.715$\pm$0.120 / 0.372$\pm$0.100 0.701$\pm$0.110 / 0.372$\pm$0.103 0.797$\pm$0.121 / 0.379$\pm$0.112 0.813$\pm$0.113 / 0.502$\pm$0.154 0.812$\pm$0.118 / 0.325$\pm$0.144 0.810$\pm$0.090 / 0.428$\pm$0.165
@@
Match Reranker. We first examine the impact of using the bipartite strategy (Section 2.2). Figure 10 shows the increase/decrease in accuracy from Magneto-zs, which directly returns the retriever matches, to Magneto-zs-bp. Note that the bipartite approach can significantly improve Recall@GT. All datasets benefit from the technique at some point, except Magellan, whose scores are already very high. The absolute increase can reach nearly 0.6. The improvements for MRR are lower but still sizable for ChEMBL and OpenData.
We also evaluate: (1) improvements of the LLM-based reranker over the bipartite method, (2) performance variations with more candidates sent to the LLM, and (3) outcomes when bypassing the SLM retrieval stage. We focus on the GDC dataset, noted for its high column count, using zero-shot settings for the retriever. Figure 11 presents accuracy and runtime for different numbers of candidates ($k$) processed by GPT-4o-mini in Magneto-zs-llm, comparing it against the use of LLaMA3.3-70B (Llama) and Magneto-zs-bp (BP).
@@
Figure 11. Ablation of candidate counts ($k=3$ to $20$) sent to GPT-4o-mini in Magneto-zs-llm. We also compare baselines that use Magneto-zs-bp (BP) for reranking, and use only an LLM – and LLaMA3.3-70B (Llama) and GPT-4o-mini (All Schema). Runtime shown on a logarithmic scale.
The LLM-based re-ranking approach improves over the bipartite. For example, at $k=5$, we observe a 6.7% improvement in MRR over the bipartite baseline, rising from 0.731 to 0.780. Recall@GT shows even more promising gains, from 0.375 to 0.475—a 26.7% increase. Increasing $k$ enhances MRR, although non-forwarded candidates score lower, reducing Recall@GT which considers all pairs.
When all schemas sent to the LLM reranker, both GPT-4o-mini (All Schema) and LLaMA3.3-70B (Llama), the performance is lower than all Magneto variations. Furthermore, runtime increases with column count and is substantial when reranking all schemas, reaching over 6,000 seconds per table for GPT-4o due to LLM API latency. In contrast, when the SLM-based retriever is used, the runtime remains practical. We used $k=20$ for our other experiments, as it balances good MRR, Recall@GT, and runtime across GDC and other datasets. These results demonstrate the effectiveness of combining SLMs and LLMs for schema matching.